OpenStudy (anonymous):

Solve S = 3f - 12 for f

OpenStudy (jhannybean):

Now just like our last problem,what an you do to keep your f's on one side and your constants on the other?

OpenStudy (jhannybean):

can*

OpenStudy (jhannybean):

Think of "S" as another number, and your "f" as your variable x.We're solving for x. X has to stay on one side while everything else moves over to the right side, yes?

OpenStudy (anonymous):

a) f= s+3 over 12 b) f= s=12 over 3 c) f= s+12 d) f= s-12

OpenStudy (anonymous):

add 12 to both sides, then divide by 3

OpenStudy (jhannybean):

<_<...

OpenStudy (anonymous):

add 12 to both sides then divide by 3? divide what by 3

OpenStudy (anonymous):

thank you got it(:

OpenStudy (anonymous):

let me show you... s = 3f - 12 (add 12 to both sides) s + 12 = 3f - 12 + 12 s + 12 = 3f (s + 12)/3 = f